### v3

# Fully tensorial approach to hypercomplex neural networks

Abstract
Fully tensorial theory of hypercomplex neural networks is given. It allows neural networks to use arithmetic based on arbitrary algebras. The key point is to observe that algebra multiplication can be represented as a rank three tensor and use this tensor in every algebraic operation. This approach is attractive for neural network libraries that support effective tensorial operations. It agrees with previous implementations for four-dimensional algebras.
keywords:
Tensor, Hypercomplex-valued Neural Network, Hypercomplex Algebra, TensorFlow, PyTorch
journal: ArXiv
[…]
[inst1]organization=Faculty of Computer Science and Telecommunications, Cracow University of Technology, addressline=Warszawska 24, city=Cracow, postcode=PL31-155, country=Poland
1 Introduction
The fast progress in applications of Artificial Neural Networks (NN) promotes new directions of research and generalizations. This involves advanced mathematical concepts such as group theory [24], differential geometry [6], [7], or topological methods in data analysis [8].
The core of NN implementations lies in linear algebra usage. In most popular programming libraries (TensorFlow [1], PyTorch [19]), the common architecture is the feed-forward NN, which is based on a stack of layers where the data passes between them unidirectionally. Optimized tensorial (multidimensional arrays) operations realize the flow of the data.
There are different algebraic extensions. One of these paths follows Algebraic Neural Networks [18], where the additional endomorphism operations on data are performed. The other algebra-geometry direction is the neural networks based on Geometric Algebra/Clifford algebra [5], [21], starting from the most prominent example of Quaternions [11], [14]. Recently, Parametrized Hypercomplex Neural Networks were invented [13] for convolutional layers. They can learn optimal hypercomplex algebra adjusted to data exploring optimized Kronecker product. However, some applications need hyperalgebra, or even more general algebra parameters, as (fixed) hyperparameters. In such a way we can optimize algebra structure at the metalevel.
In this paper, we discuss implementation in which we change classical real algebra computations into various hypercomplex or even more general algebra computations by including multiplication in the algebra as a tensorial operation. Hypercomplex NN approach is not new, e.g., [3], [11]. However, interest in this direction is revived due to better complexity properties in such areas as image processing [23] or time series analysis [15]. In these contributions the Open Source code [23] for specific four-dimensional hypercomplex algebras were used.
The implementation explained in this article significantly expands the ideas from [23], [11] for arbitrary algebras, including hypercomplex ones. The algorithms described here agree with the NN presented in [23], [11] when limited to 4-dimensional hypercomplex algebras. The 4-dimensional hyperalgebras are useful in encoding RGB color data in images. Moreover, implementation of [23] was obtained by constructing an additional multiplication matrix from the multiplication table for the hypercomplex algebra. This is an additional step in setting up a neural network. Our approach permits us to omit this complexity and generalize to arbitrary algebras since we input only the multiplication table of the algebra as it can be found in standard references, e.g., [2]. Moreover, the limitation to 4-dimensional algebras can be relaxed. This is important for processing data that naturally encodes in $n$-tuples ($n>0$, integer) that can be encoded as a single element of an $n$-dimensional algebra. This is a crucial contribution from the theoretical treatment of the general algebraic approach to hypercomplex neural networks.
The main contribution of this paper is the following:
  • 1.
    summarize basic concepts of hyperalgebras and more general algebras, and prove that the algebra multiplication can be expressed as a third-rank tensor,
  • 2.
    provide a general algorithm for computations within the hypercomplex dense layer that generalizes ideas of [23], [11] for arbitrary algebras of arbitrary algebra dimensions,
  • 3.
    provide general algorithms for 1-, 2-, and 3-dimensional hypercomplex convolutional layer computations that generalizes ideas of [23], [11] for arbitrary algebras of any algebra dimensions.
The paper is organized as follows: after recalling the general algebraic notions in the next section, the algorithms for general architectures of dense and convolutional hypercomplex NN are presented. Finally, in Section 5, we present some conclusions and future directions for work.
2 Theory
This section provides an overview of the mathematical theory behind the operations used in implementing hypercomplex neural networks. This is a tenet of methods used in this paper. A tensor is a classical multilinear algebra notion explained in detail in standard references, e.g., [2]. We rephrase it here to prepare the background for explaining how algebra multiplication can be expressed as a third-rank tensor.
2.1 Tensors
The primary object used in NN implementations is a ’tensor’. In typical NN libraries such as TensorFlow and PyTorch tensors are synonyms of multidimensional matrix. In algebra, tensors, apart from being a mathematical object that can be represented as multidimensional arrays after selecting a base of a vector space, have many interesting and valuable properties. It relies on the tensor product.
We will consider vector spaces: $V$ and its dual $V^{*}$ over a field $\mathbb{F}$ that can represent, e.g., real numbers. For a fixed base $\{e_{i}\}$ of $V$ the dual space has a base $\{e^{j}\}$, where $e^{i}(e_{j})=\delta^{i}_{j}$, where $\delta^{i}_{j}$ is the Kronecker delta defined as follows $\delta^{i}_{j}=\left\{\begin{array}[]{cc}1&i=j\\ 0&i\neq j\end{array}\right.$.
Definition
The tensor product of two vector spaces $V_{1}$ and $V_{2}$ over the field $\mathbb{F}$ is the vector space denoted by $V_{1}\otimes V_{2}$ and defined as a quotient space $V_{1}\times V_{2}/L$, where the subspace $L\subset V_{1}\times V_{2}$ is spanned by
$\begin{array}[]{c}(v+w,x)-(v,x)-(w,x),\\ (v,x+y)-(v,x)-(v,y),\\ (\lambda v,x)-\lambda(v,x),\\ (v,\lambda x)-\lambda(v,x),\end{array}$ (1)
where $v,w\in V_{1}$, $x,y\in V_{2}$, $\lambda\in\mathbb{F}$. The equivalence class of $(v,x)$ in $V_{1}\otimes V_{2}$ is denoted by $v\otimes x$.
From practical point of view, for $v+\lambda w\in V_{1}$ and $x+y\in V_{2}$, $\lambda\in\mathbb{F}$, then $(v+\lambda w)\otimes(x+y)=v\otimes x+v\otimes y+\lambda w\otimes x+\lambda w\otimes y$ belongs to $V_{1}\otimes V_{2}$.
By induction, the tensor product can be defined for $k$ vector spaces $\{V_{i}\}_{i=1}^{k}$ and is denoted by $V_{1}\otimes\ldots\otimes V_{k}$ or shortening notation as $\otimes_{i=1}^{k}V_{i}$.
The tensor product can also be defined for dual spaces $V_{i}^{*}$ of a vector spaces $V_{i}$ for $i\in\{1,\ldots k\}$, $0<k<\infty$, and we can define mixed tensor product made from vector spaces and their duals.
The tensor product of a vector space is a vector space, so we can define a base, e.g., if the base of $V$ is $\{e_{i}\}_{i=1}^{n_{1}}$ and $W$ is $\{f_{i}\}_{i=1}^{n_{2}}$, then the base of $V\otimes W$ is $\{e_{i}\otimes f_{j}\}_{(i,j)=(1,1)}^{(n_{1},n_{2})}$.
The tensor product of vector spaces is useful in decomposing any multilinear mapping. It is expressed in the so-called universal factorization theorem for tensor product [2]. For a bilinear mapping $F:V\times W\rightarrow\mathbb{F}$ of vector spaces $V,W$, it states that $F$ can be uniquely factorized by a new mapping $f:V\otimes W\rightarrow\mathbb{F}$, which is expressed by the commutative diagram in Fig. 1.
[…]
Figure 1: Unique factorization property of the tensor product means that $F=f\circ\otimes$.
Then, the map $f$ is called the tensor. This theorem can be generalized to any multilinear mapping.
Definition
A multilinear mapping $F:\underbrace{V\times\ldots\times V}_{p}\times\underbrace{V^{*}\times\ldots\times V^{*}}_{q}\rightarrow\mathbb{F}$ of a vector space $V$ and its dual $V^{*}$, can be factorized by the tensor $f:\otimes_{i=1}^{p}V\otimes\otimes_{j=1}^{q}V^{*}\rightarrow\mathbb{F}$ by $F=f\circ\otimes$.
For fixed bases $\{e_{i}\}$ of $V$ and its dual $\{e^{j}\}$ for $V^{*}$ we can decompose $f$ as [^1] Here we use Einstein summation convention: repeated bottom and top index indicate the summation over the whole range.
$f=f_{i_{1}\ldots i_{p}}^{j_{1}\ldots j_{q}}e^{i_{1}}\otimes\ldots\otimes e^{i_{p}}\otimes e_{j_{1}}\otimes\ldots\otimes e_{j_{q}},$ (2)
where numbers $f_{i_{1}\ldots i_{p}}^{j_{1}\ldots j_{q}}$ are components of the tensor $f$. The tensor $f$ is called $p$-covariant and $q$-contravariant or $(p,q)$ tensor.
The tensor components can be arranged in a multidimensional array and are called ’tensor’ in packages such as TensorFlow or PyTorch. We can also adopt the concept of a tensor’s shape, which is a tuple containing the range of all indices. The nomenclature in these libraries names the axis each index, which represents a vector space in the tensor product.
Example
In the bases $\{e_{i}\}$ of $V$ and $\{f_{j}\}$ of $W$ the bilinear mapping $F:V\times W\rightarrow\mathbb{R}$ factorizes through tensor product space as
$F(e_{i},f_{j})=f(e_{i}\otimes f_{j})=F_{ij},$ (3)
where $F_{ij}\in\mathbb{R}$ for all $i,j$, are the coefficients of a numerical matrix that is a representation of the bilinear mapping (tensor) $F$ in the fixed base of $V$ and $W$. The matrix collects the components of the tensor in a fixed tensor base. We can therefore write tensor in the base $\{e^{i}\}$ of $V^{*}$ and $\{f^{j}\}$ of $W^{*}$, i.e., $e^{i}(e_{j})=\delta^{i}_{j}=f^{i}(f_{j})$, as
$f=F_{ij}e^{i}\otimes f^{j}.$ (4)
We have for two vectors $u=u^{i}e_{i}$, $x=x^{i}f_{i}$
$F(u,x)=f(u\otimes x)=F_{ij}e^{i}\otimes f^{j}(u\otimes x)=F_{ij}e^{i}(u)f^{j}(x)=F_{ij}u^{i}x^{j}.$ (5)
The tensor $f$ has the shape $(dim(V),dim(W))$.
Let us notice that the matrix $[F_{ij}]$ is implemented in TensorFlow and PyTorch library as a tensor class. The critical difference is that mathematical tensors have specific properties of transformations under the change of basis in the underlying vector spaces. The libraries implement the tensors as a multidimensional matrix of numbers. Moreover, the libraries usually do not keep the upper (contravariant) or lower (covariant) position of indices.
The example can be extended to the tensor product of multiple vector spaces and their duals.
Example
The linear operator $F:V\rightarrow W$ can be written as a mapping $f\in V^{*}\otimes W\rightarrow\mathbb{R}$ with partial application of a vector. It can be written in the base $\{e^{i}\}$ of $V^{*}$ and $\{f_{i}\}$ of $W$ as $f=F_{i}^{j}e^{i}\otimes f_{j}$, where $F_{i}^{j}$ is the matrix of $F$ in the mentioned bases of $V$ and $W$. The vector space of linear operators is denoted as $L(V,W)$. One can prove that there is a bijection between linear operators and mixed rank $2$ tensors.
For tensors we also often use abstract index notation where we provide only components of the tensor, e.g., $A_{ij}$, understanding them not as fixed base numerical values but as a full tensor $A_{ij}e^{i}\otimes e^{j}$ (cf. [20]).
Since the tensor product is a functor in the category of linear spaces [2], therefore, for a linear mapping $f:V\rightarrow W$, we can define the extension of the mapping for a tensor product space $\otimes_{i=1}^{n}f:\otimes_{i=1}^{n}V\rightarrow\otimes_{i=1}^{n}W$ by acting on product as
$\otimes f(v_{1}\otimes\ldots\otimes v_{n}):=fv_{1}\otimes\ldots\otimes fv_{n},$ (6)
and extending by linearity for all combinations. This is similar rule as for extending linear operators to the Cartesian product of vector spaces. This results from the fact that the Cartesian product is also a functor, just as a tensor product is.
We can define a few linear algebra operations realized in tensor libraries.
  • 1.
    Broadcasting: it is defined for a linear operator $A:V\rightarrow W$ to be a multilinear extension to the Cartesian product of spaces
    $b_{1}^{n}:L(V,W)\rightarrow L(\times_{i=1}^{n}V,\times_{i=1}^{n}W),$ (7)
    that is $b_{1}^{n}(A)(v_{1},\ldots,v_{n})=(Av_{1},\ldots,Av_{n})$. Similar broadcasting is realized for the tensor product. Broadcasting relies on the functionality of the Cartesian product and tensor products.
  • 2.
    Transposition/Permutation: The transposition of two components relies on the following fact: there is isomorphism $\tau:V\otimes W\rightarrow W\otimes V$ that simply reverses the order of factors $\tau(v\otimes w)=w\otimes v$. We can extend it to arbitrary permutation of $n$ numbers, that is, a bijective mapping: $p:\{1,\ldots,n\}\rightarrow\{1,\ldots,n\}$. We define
    $\tau_{p}(v_{1}\otimes\ldots\otimes v_{n})=v_{p(1)}\otimes\ldots\otimes v_{p(n)}$ (8)
    We can define a similar operation for the Cartesian product.
    In the abstract index notation, we have $\tau_{p}A_{i_{1}\ldots i_{k}}=A_{i_{p(1)}\ldots i_{p(k)}}$.
  • 3.
    Reshaping: For a pair of indices $i,j$ of the range $0\leq R(i)<\infty$ and $0\leq R(j)<\infty$, where $R$ denotes the maximal value of its argument, which is an index. We can define a reshape operation $Rs$, which is the new index $Rs(i,j)$ that value depends on the values of $i,j$ given by the function
    $Rs(i,j)=iR(j)+j.$ (9)
    We can extend the operation for the pair $(p,p+1)$ of neighbor indices of a tensor, using abstract index notation, by $Rs_{p}A_{i_{1}\ldots i_{p}i_{p+1}\ldots i_{k}}=A_{i_{1}\ldots Rs(i_{p},i_{p+1})\ldots i_{k}}$, where $p+1\leq k$. This operation changes only the way of indexing; however, it is useful in applications.
  • 4.
    Contraction: For a two tensors $A\in\ldots\otimes\underbrace{V}_{p}\otimes\ldots\rightarrow\mathbb{F}$ and $B\in\ldots\otimes\underbrace{V^{*}}_{q}\otimes\ldots\rightarrow\mathbb{F}$, and a fixed base $\{e_{i}\}$ of $V$ and $\{e^{j}\}$ of $V^{*}$, the contraction of indices $p$ (related to $V$ in $A$) and $q$ (related to $V^{*}$ in $B$) is (note implicit sum):
    $C_{p,q}(A,B)=A(\ldots\underbrace{e_{i}}_{p}\ldots)\otimes B(\ldots\underbrace{e^{i}}_{q}\ldots),$ (10)
    or in abstract index notation
    $C_{p,q}(A,B)=A_{\ldots\underbrace{i}_{p}\ldots}\otimes B^{\ldots\overbrace{i}^{q}\ldots}.$ (11)
    The contraction can also be defined for a single tensor in the same way, e.g., in abstract index notation for a single tensor $T$, ones get $C_{p,q}T_{\ldots i_{p}\ldots}^{\ldots i_{q}\ldots}=T_{\ldots\underbrace{i}_{p}\ldots}^{\ldots\overbrace{i}^{q}\dots}$, where implicit summation was applied.
  • 5.
    Concatenation: joins the tensors $\{A^{(i)}_{i_{1}\ldots i_{l}}\}_{i=1}^{n}$ of the same shape along given dimension $j$, i.e.,
    $K_{j}(\{A^{(i)}\}_{i=1}^{n})=A_{i_{1}\ldots i_{j-1}k_{j}i_{j}\ldots i_{l}}=A^{(k_{j})}_{i_{1}\ldots i_{j-1}i_{j}\ldots i_{l}}.$ (12)
2.2 (Hypercomplex) Algebras
In this part we introduce mathematical concepts related to hypercomplex and general algebras. We start from
Definition
Algebra over a field $\mathbb{F}$ is a vector space $V$ equipped with a product - a binary operation __$\_\cdot\_:V\times V\rightarrow V$ with the following properties:
  • 1.
    $(x+y)\cdot z=x\cdot z+y\cdot z$,
  • 2.
    $z\cdot(x+y)=z\cdot x+z\cdot y$,
  • 3.
    $(\alpha x)\cdot(\beta y)=\alpha\beta(x\cdot y)$,
for $x,y,z\in V$ and $\alpha,\beta\in\mathbb{F}$.
The algebra is commutative if $x\cdot y=y\cdot x$ for $x,y\in V$.
We can define multiplication in the algebra by defining it on the base $\{e_{i}\}_{i=1}^{dim(V)}$ of underlying vector space $V=span(\{e_{i}\})$.
Example
For real numbers $\mathbb{R}$ we have $V=span(\{e_{0}\})$, $\mathbb{F}=\mathbb{R}$ with $e_{0}\cdot e_{0}=e_{0}$. We generally associate $e_{0}$ with $1\in\mathbb{R}$.
Example
Complex numbers $\mathbb{C}$ can be obtained from commutative algebra with $V=span(\{e_{0},e_{1}\})$, $F=\mathbb{R}$ and $e_{0}\cdot e_{0}=e_{0}$, $e_{0}\cdot e_{1}=e_{1}$ and $e_{1}\cdot e_{1}=-e_{0}$. We associate $e_{0}$ with $1$, and $e_{1}$ with the imaginary unit $i$.
By convention we will always assume that the neutral element of the algebra multiplication, denoted by $1$, is $e_{0}$.
Efficient use of algebras in computations based on tensors (in TensorFlow and PyTorch) relies on converting the product within the algebra into a tensor operation.
Treating algebra as a vector space with the additional structure of vector multiplication [^2] It can be defined using a functor that casts the category of algebras into the category of linear spaces. we have the following Theorem, which is essential to the rest of the paper.
Theorem
For an algebra $V$ over $F$ the product can be defined as a tensor $A:V^{*}\otimes V^{*}\otimes V\rightarrow\mathbb{F}$. Selecting the base of $V=span(\{e_{i}\}_{i=0}^{n})$, $n=dim(V)<\infty$, and the dual base $V^{*}=span(\{e^{i}\}_{i=1}^{n})$ with $e^{i}(e_{j})=\delta_{j}^{i}$, the product is defined by the tensor
$A=A_{ij}^{k}e^{i}\otimes e^{j}\otimes e_{k}.$ (13)
The multiplication is defined by partial application of vectors $x=x^{i}e_{i},y=y^{i}e_{i}\in V$
$x\cdot y:=A(x,y)=A^{k}_{ij}x^{i}y^{j}e_{k}.$ (14)
The multiplication table entry is presented in [15].
$\begin{array}[]{c|c}&e_{j}\\ \hline\cr e_{i}&A_{ij}^{k}e_{k}\end{array}$ (15)
Proof.
For the proof, we must check if the multiplication defined by $A$ fulfills the axioms of multiplication $x\cdot y$ for an algebra.
By linearity of tensor $(x+y)\cdot z=A(x+y,z)=A(x,z)+A(y,z)=x\cdot z+y\cdot z$, and $x\cdot(y+z)=A(x,y+z)=A(x,y)+A(x,z)=x\cdot y+x\cdot z$, for $x,y,z\in V$.
Moreover, if $\alpha,\beta\in\mathbb{F}$ and $x,y\in V$ we have by linearity of tensor $(\alpha x)\cdot(\beta y)=A(\alpha x,\beta y)=\alpha\beta A(x,y)=\alpha\beta(x\cdot y)$, as required. ∎
Corollary When the algebra is commutative, then $A(x,y)=A(y,x)$ or $A_{ij}^{k}=A_{ji}^{k}$.
Proof.
Since $x\cdot y=y\cdot x$, so $A(x,y)=A(y,x)$, which results in symmetry of tensor indices $A_{ij}^{k}=A_{ji}^{k}$, as required. ∎
The tensor coefficients (abstract index notation), $A_{ij}^{k}$, play the same role as structure constants for a group [10]. These coefficients in a fixed base can be represented as a multidimensional matrix (called tensors in TensorFlow and PyTorch).
Example For complex numbers $\mathbb{C}$, $V$ has two dimensional base $\{e_{0},e_{1}\}$ with multiplication matrix given in [16].
$\begin{array}[]{c|c|c}\cdot&e_{0}&e_{1}\\ \hline\cr e_{0}&e_{0}&e_{1}\\ \hline\cr e_{1}&e_{1}&-e_{0}\\ \end{array}$ (16)
Then the multiplication can be written as the following tensor
$A_{\mathbb{C}}=e^{0}\otimes e^{0}\otimes e_{0}+e^{0}\otimes e^{1}\otimes e_{1}+e^{1}\otimes e^{0}\otimes e_{1}-e^{1}\otimes e^{1}\otimes e_{0}.$ (17)
3 Neural Networks Architectures
In this section, we provide mathematical details of implementing hypercomplex dense and convolutional neural networks.
We do not distinguish co- and contravariant indices in the description, and we write them in subscripts. Moreover, the tensor is treated as a multidimensional array, with indices starting from 0. This is a standard convention in the tensor libraries such as TensorFlow and PyTorch.
We assume that the algebra multiplication structure constant is fixed and stored in the tensor (in abstract index notation) $A=A_{i_{a}j_{a}k_{a}}$ (cf. Subsection 2.2).
We want to stress out that the proposed algorithms are generalizations of algorithms from [23], [11] from $4$-dimensional algebras to arbitrary dimensional algebras using tensor operations that are optimized for tensor operations from packages like TensorFlow or PyTorch. They are implemented in Hypercomplex Keras Python library [25] and described elsewhere [17]. The implementation was constructed to agree for $4$-dimensional algebras with the results from [23].
3.1 Hypercomplex dense layer
We start with a description of the dense hypercomplex layer. It is a general-purpose layer that operates on the data with additional dimensionality/axis, which is a multiple of algebra dimensions. We assume that the input data are of dimension $b\times\underbrace{al\times in}_{n}$, where $b$ is the batch size, $al$ the algebra size, and $in$ the positive integer multiplier. The last two numbers determine the input data size $n$. The input tensor in abstract index notation is $X=X_{i_{b}Rs(i_{al},i_{in})}$, where $i_{b}$ batch index, $i_{al}$ is the algebra index, $i_{in}$ is the multiplicity index of algebra dimension. Moreover, we use learning parameters (weights/kernel) $K=K_{i_{al}i_{in}i_{u}}$, where $i_{u}$ is the index over units/neurons. The bias $b=b_{Rs(i_{al},i_{u})}$ is used if needed. Kernel and bias are usually initialized with numbers taken from specific distributions (cf. [12]).
The algorithm of the hypercomplex dense layer is provided in Algorithm 1. We offer both tensorial and abstract index notations (AIN). Since indices are in the subscript, so we write the summation sign explicitly. We need two flags bias - if bias is included and activation - if activation function $\sigma$ is used.
Algorithm 1 Hypercomplex dense NN
$X$, $A$, $\sigma$, bias, activation
$K$, $b$ initialized
$W\leftarrow C_{1,0}(A,K)$ [AIN: $W_{i_{a}k_{a}i_{in}i_{u}}\leftarrow\sum_{j}A_{i_{a}jk_{a}}K_{ji_{in}i_{u}}$]
$W\leftarrow\tau_{p=(0\rightarrow 0,1\rightarrow 2,2\rightarrow 1,3\rightarrow 3)}(W)$ [AIN: $W_{i_{a}i_{in}k_{a}i_{u}}\leftarrow W_{p(i_{a}k_{a}i_{in}i_{u})}$]
$W\leftarrow Rs_{1}\circ Rs_{0}(W)$ [AIN: $W_{i_{1}i_{2}}\leftarrow W_{Rs(i_{a},i_{in})Rs(k_{a},i_{u})}$]
$Output\leftarrow C_{1,0}(X,W)$ [AIN: $Output_{i_{b}i_{2}}\leftarrow\sum_{k}X_{i_{b}k=Rs(i_{al},i_{in})}W_{ki_{2}}$]
if bias is True then
     $Output\leftarrow Output+b$
end if
if activation is True then
     $Output\leftarrow b(\sigma)(Output)$
end if
The implementation using Keras with TensorFlow library implementation and PyTorch library implementations are described in [17] and implemented in Hypercomplex Keras library [25] in the HyperDense class. As it was emphasized, for $4$-diemnsional algebras the output agrees with the results of [23].
3.2 Hypercomplex convolutional layer
In this part the hypercomplex convolutional neural network will be described. We present general $k$-dimensional ($k=1,2,3$) layers. They differ by the shape of the input data and kernel size.
The additional ’image channels’ of the data (typically $3$-dimensional subarray of RGB colors) can be packed as an element of algebra. For instance, the two-dimensional image can be decomposed as a matrix of four-dimensional algebra elements, i.e., color channels plus alpha value, making a single pixel an element of four-dimensional algebra.
The general idea of the NN operation is to use algebra multiplication matrix tensor $A$ to separate the coefficients of the algebra base component and then apply the traditional convolution for each algebra component. This idea is an extension of the results of [23] for four-dimensional algebras, and in particular, of [11] for Quaternions. In [23], the split of components of algebra was implemented at the level of a specially tailored multiplication matrix, so it was not general.
The dimension of the input data $X=X_{i_{b}i_{1}\ldots i_{k}Rs(i_{al},i_{in})}$ of dimension $b\times n_{1}\times\ldots\times n_{k}\times\underbrace{al\times in}_{n}$, where $b$ is the batch size, $n_{i}$, $i\in\{1,\ldots,k\}$ the size of data sample in each dimension (e.g., for $k=2$ the raster coordinates $i_{1}$, $i_{2}$), $al$ the algebra dimension, and $in$ the positive multiplier [^3] One can see that the size of the data sample in each dimension (axes) stands right after the batch size, which is used in TensorFlow convention. For PyTorch, this multiindex is placed at the end. We will not provide an implementation for the PyTorch convention since it differs by a few permutations.. The kernel size is $al\times L_{1}\times\ldots L_{k}\times in\times F$, where: $L_{i}$, $i\in\{1,\ldots,k\}$ is the kernel dimension for each axis, and $F$ the number of filters. We therefore have the kernel $K=K_{i_{al}i_{l_{1}}\ldots i_{l_{k}}i_{in}i_{f}}$. The flag $bias$ indicates the bias $b$ usage. If it is used, it has dimension $al\times F$. Kernel and bias can be initialized by arbitrary distributions; see [12]. We use standard $k$-dimensional convolution $convkD(X,K,strides,padding)$ for convoluting algebra components, there are standard optimized convolution operations [16]. The algorithm is presented in the Algorithm 2.
Algorithm 2 Hypercomplex $k$-dimensional convolutional NN
$X$, $A$, $\sigma$, bias, activation
$K$, $b$ initialized
$W\leftarrow C_{1,0}(A,K)$ [AIN: $W_{i_{a}k_{a}i_{l_{1}}\ldots l_{l_{k}}i_{in}i_{f}}\leftarrow\sum_{j}A_{i_{a}jk_{a}}K_{ji_{l_{1}}\ldots i_{l_{k}}i_{in}i_{f}}$]
$W\leftarrow\tau_{p=(0\rightarrow k+2)}(W)$ [AIN: $W_{k_{a}i_{l1}\ldots l_{lk}i_{a}i_{in}i_{f}}\leftarrow W_{p(i_{a}k_{a}i_{l_{1}}\ldots l_{l_{k}}i_{in}i_{f})}$]
$W\leftarrow Rs_{k+2}(W)$ [AIN: $W_{k_{a}i_{l_{1}}\ldots l_{l_{k}}ji_{f}}\leftarrow W_{k_{a}i_{l_{1}}\ldots l_{l_{k}}\underbrace{Rs(i_{a},i_{in})}_{=j}i_{f}}$]
temp = []
for $i\in\{0,\ldots,al\}$ do temp = temp + [$convkD(X,W_{i\ldots})$]
end for
$Output_{i_{b}i_{l1}\ldots i_{lk}h}\leftarrow K_{k+1}(temp)$
if bias is True then
     $Output\leftarrow Output+b$
end if
if activation is True then
     $Output\leftarrow b(\sigma)(Output)$
end if
The algorithm for different $k\in\{1,2,3\}$ is implemented in the classes HyperConv1D, HyperConv2D, and HyperConv3D of Hypercomplex Keras library [25] and described in [17]. It agrees and extends the results of [23] from $4$-dimensional algebra to arbitrary dimensional algebras.
4 Tensor version of universal approximation theorem for single layer of hypercomplex perceptrons
In this section we give the proof of Universal Approximation Theorem in the tensor framework outlined above. It presents how this approach provides more streamlined approach. We will follow [3]. For alternative vector-matrix approach see [22], which is also based on [3].
In this section we assume that the algebra $(V,+,\cdot)$ is non-degenerate, i.e.,
Definition 1.
An algebra $(V,+,\cdot)$ with multiplication given by the tensor $A$ is non-degenerate if for $x\in V$
$\begin{array}[]{cccc}A(x,y)=0,&\forall y\in V&\Rightarrow&x=0,\end{array}$ (18)
and
$\begin{array}[]{cccc}A(y,x)=0,&\forall y\in V&\Rightarrow&x=0.\end{array}$ (19)
The crucial part is played by the Riesz’s representation theorem that in standard form is [4]
Lemma 1 (Riesz’s representation lemma).
For a Hilbert space $H$ over field $\mathbb{F}$ (e.g., $\mathbb{R}$ or $\mathbb{C}$) with inner product <_|_$<\_|\_>:X\times X\rightarrow\mathbb{F}$ every continous linear functional $\phi\in H^{*}$ there is a vector $v_{\phi}\in H$ such that
$\phi(x)=<v_{\phi}|x>,$ (20)
for all $x\in H$.
When we take $H=\mathbb{R}^{N}$, for integer $N>0$, then $<x|y>:=x^{T}y=\sum_{i=1}^{n}x_{i}y_{i}$. Moreover, if $dim(H)<\infty$, then the linear functional is automatically continuous.
In order to amend the lemma for vectors that entries are hypercomplex numbers from $V$, we define projection on components of hypercomplex numbers $\pi_{k}(\sum_{i=0}^{n}x^{i}e_{i})=x^{k}$ for $0\leq k\leq n$. We then prove the generalization of Lemma 5.5.1 from [3]
Lemma 2 (Riesz’s representation lemma for nondegenerate algebra).
Let $\phi:V^{N}\rightarrow\mathbb{R}$ be a linear functional, where $V$ is a nondegenerate algebra. Then there exists, uniquely determined for $\phi$, the set of vectors $\{v_{i}\}_{i=0}^{n}$ for $v_{i}\in V^{N}$ such that for every $x\in V^{N}$ we have
$\phi(x)=\pi_{i}(<v_{i}|x>)$ (21)
for all $i\in\{0,\ldots,n\}$, and $<v|x>=\sum_{k=1}^{N}v_{k}\cdot x_{k}=v^{T}x$.
Proof.
We fix $0\leq i\leq n$ and we have $\phi(x)=\pi_{i}(<v_{i}|x>)$. Define a morphism $\Phi:V^{n}\rightarrow(V^{n})^{*}$ such that $\Phi(v_{i})=\phi=\pi_{i}\circ<v_{i}|\_>$. We prove that $\Phi$ defines an isomorphism.
One can note that $\Phi$ is linear and the image and preimage have the same dimension. Therefore, in order to prove that $\Phi$ is an isomorphism, we prove that $\Phi$ is injective, i.e., if $\Phi(v_{i})(x)=\pi_{i}(<v_{i}|x>)=\phi(x)=0$ for all $x\in V^{N}$, then $v_{i}=0$.
Selecting base vectors $E_{k}=[0,\ldots,0,\underbrace{1}_{k},0,\ldots,0]\in V^{N}$ for $1\leq k\leq N$ we define vectors $y_{pq}=e_{p}E_{q}$ where index $0\leq p\leq n$ runs over algebra base of $V$, and $q$ runs over components of $V^{N}$. The vectors $y_{pq}$ form a base of vectors with components from the algebra $V$. Assuming that
$0=\phi(y_{pq})=\pi_{i}(<v_{i}|y_{pq}>),$ (22)
for all $p$ and $q$, and taking into account that $V$ algebra is nondegenerate, we have that all components of $v_{i}$ must vanish, so $v_{i}=0$. As $i$ is arbitrary, this is true for all $0\leq i\leq n$ so the lemma is proved. ∎
We can now prove the Universal Approximation Theorem for hypercomplex-valued vectors. For the proof based on vector-matrix notation see [22].
Theorem 1.
Let $X\subset V^{N}$ be a compact subset and let $g:X\rightarrow V^{N}$ be a continuous function. Then for all $\varepsilon>0$ there exist some coefficients:
$\alpha_{1},\ldots,\alpha_{p}\in\mathbb{R},$ (23)
some vectors:
$v_{1},\ldots,v_{p}\in V^{N}$ (24)
and some elements of algebra:
$\Theta_{1},\ldots,\Theta_{p}\in V,$ (25)
such that:
$sup_{x\in X}||g(x)-\sum_{k=1}^{p}\alpha_{k}f(<v_{k}|x>+\Theta_{k})||<\varepsilon,$ (26)
for $f=\sum_{i=1}^{n}\sigma\circ\pi_{i}$, where $\sigma$ is the sigmoid function, and the norm is a vector space norm on an algebra $V$ treated as a vector space.
In other words, the set of functions $S=\sum_{k=1}^{p}\alpha_{k}f(<v_{k}|x>+\Theta_{k})$ for parameters: $\alpha_{k}$, $v_{k}$ and $\Theta_{k}$ are dense in $C^{0}(X,V)$ continuous functions from $X$ to $v$ with the supremum norm over components of $V$ treated as a vector space.
Proof.
The general idea of the proof is as follows:
  1. 1.
    Decompose the function $g$ into components of the algebra $V$. The component is a real-valued function.
  2. 2.
    Use classical Universal Approximation Theorem for the component. This hides all Functional Analysis parts of the proof to already proven the classical theorem [9].
  3. 3.
    Compose components approximations into the approximation of whole function using supremum norm for estimate approximation quality.
To begin with, consider a function $g:X\rightarrow V$ that can be decomposed into components of algebra $V$ as
$g=\sum_{i=0}^{n}e_{i}\otimes\pi_{i}\circ g.$ (27)
Then each component $\pi_{i}\circ g\in C^{0}(X,\mathbb{R})$ can be approximated by the Cybenko’s Universal Approximation Theorem [9]: there are real coefficients $a_{1}^{i},\ldots,a_{p}^{i}$, real-valued linear functionals $\phi_{1}^{i},\ldots,\phi_{p}^{i}$ on $\mathbb{R}^{nN}=V^{n}$ and some real parameters $\theta_{1}^{i},\ldots,\theta_{p}^{i}$, where the index depends on the component $i$, i.e., $p=p(i)$, such that
$\left|\pi_{i}(g(x))-\sum_{k=1}^{p}a_{k}^{i}\sigma(\phi_{k}^{i}(x)+\theta_{k}^{i})\right|<\varepsilon$ (28)
for all $x\in X$.
From the Lemma 2 we have that $\phi_{k}^{i}=\pi_{i}\circ<v_{k}^{i}|\_>$ for a fixed vector $v_{k}^{i}\in V^{N}$.
Consider now $\Theta_{k}^{i}(\lambda^{i})=\lambda^{i}e_{0}+\ldots+\lambda^{i}e_{i-1}+\theta_{k}^{i}e_{i}+\lambda^{i}e_{i+1}+\ldots+\lambda^{i}e_{n}$. We have that
$lim_{\lambda^{i}\rightarrow-\infty}f(<v_{k}^{i}|x>+\Theta_{k}^{i}(\lambda^{i}))=\sigma(\phi_{k}^{i}(x)+\theta_{k}^{i})\otimes e_{i}$ (29)
uniformly. This results from the properties of $\sigma$ function ($lim_{\lambda\rightarrow-\infty}\sigma(\pi_{l}(<v_{k}|x>)+\lambda)=0$), which guarantees that all components apart the $i$-th one vanish in the limit. Therefore, we can select such $\lambda^{i}$ that
$sup_{x\in X}||\sum_{k=1}^{p}\alpha_{k}\sigma(\phi_{k}^{i}(x)+\theta_{k}^{i})e_{i}-\sum_{k=1}^{p}\alpha_{k}f(<v_{k}^{i}|x>+\Theta_{k}^{i}(\lambda^{i}))||<\varepsilon,$ (30)
so we get
$sup_{x\in X}||\pi_{i}(g(x))e_{i}-\sum_{k=1}^{p}\alpha_{k}f(<v_{k}^{i}|x>+\Theta_{k}^{i}(\lambda^{i}))||<2\varepsilon.$ (31)
Since it works for all components $\pi_{i}\circ g$ of $g$, so it works also for supremum of all components. Therefore $g$ can be arbitrarily approximated by $f$. The proof is complete. ∎
The above proof that mimic the proof of Theorem 5.5.1 from [3], and it delegates functional analytic parts of the proof in application of Cybenko’s Universal Approximation Theorem [9]. One can generalize the proof by replacing the sigmoid function $\sigma$ by the discriminatory sigmoidal functions [9].
5 Conclusion
We introduced the mathematical details of the implementation of dense and convolutional NN based on hypercomplex and general algebras. The critical point in this presentation is to associate algebra multiplication with rank-three tensor. Thanks to this observation, all the NN processing steps can be represented as tensorial operations.
The fully tensorial operations applied to algebra operations simplify neural network operations and allow support for fast tensorial operations in modern packages such as TensorFlow and PyTorch. Implementation of the above generalized hypercomplex NN is described elsewhere [17].
It is an extension of algorithms presented for $4$-dimensional algebras in [23] and for Quaternions in [11]. They are implemented in Hypercomplex Keras library [25] and checked to agree for $4$-dimensional algebras with the results from [23].
Supplementary information
Declarations
  • 1.
    Funding - This paper has been supported by the Polish National Agency for Academic Exchange Strategic Partnership Programme under Grant No. BPI/PST/2021/1/00031 (nawa.gov.pl).
  • 2.
    Conflict of interest/Competing interests - Not applicable
  • 3.
    Ethics approval and consent to participate - Not applicable
  • 4.
    Consent for publication - All authors agree on publication
  • 5.
    Data availability - Not applicable
  • 6.
    Materials availability - Not applicable
  • 7.
    Code availability - Not applicable
  • 8.
    Author contribution - A.N., R.K.: Conceptualization, Methodology, Investigation, Writing - Original Draft, Review, Editing. A.N.: Funding acquisition, Supervision.
Acknowledgements
This paper has been supported by the Polish National Agency for Academic Exchange Strategic Partnership Programme under Grant No. BPI/PST/2021/1/00031 (nawa.gov.pl).